# Unsupervised full-field Bayesian inference of orthotropic hyperelasticity from a single biaxial test: a myocardial case study

### v1

@@ 8 Conclusion
We presented a physics-informed stochastic variational framework that recovers the full constitutive parameter set of a nonlinear orthotropic hyperelastic model from a single biaxial stretch experiment. By exploiting intrinsic microstructural heterogeneity, or introducing simple geometric heterogeneity features when needed, the method identifies not only the dominant tensile parameters but also the shear-coupling terms that are typically weakly excited in a biaxial tensile test. Our results demonstrate that an effective balance between underlying microstructure and biaxial loading improves identifiability, and this balance can be set during slicing and tissue harvesting. Circumferential–radial slabs and rotated cuts that place both fiber and sheet directions in the loading plane enable full recovery under noise-free conditions. Under realistic displacement measurement noise, the approach remains robust and yields stresses and strain energies in close agreement with ground-truth simulations. Adding geometric heterogeneity, for example a central occlusion, systematically improves convergence and tightens posterior distributions, particularly for shear parameters. These results indicate that slicing protocols and experimental design can be used deliberately to enhance observability in full-field inverse characterization. More broadly, the framework enables localized characterization of orthotropic nonlinear biological tissues using a single-mode setup. This reduces experimental tissue manipulation burden, improves reproducibility, and supports studies where available tissue is limited.
@@ Appendix A Tissue slab geometry and loading conditions
Figure 1: Synthetic data generation: training sample dimensions and loading conditions Tissue slab geometry and applied boundary conditions for the training samples. We subject all training samples undergo to a biaxial loading protocol $\lambda_{1}:\lambda_{2}-(1:2),(1:1),(2:1)$ up to $15\%$ of the initial length of the sample. All samples are of size 10 $\times$ 10 $\times$ 1 mm3. The geometrically heterogeneous tissue samples have a circular occlusion in the center with a radius of $R=1$ mm.
Appendix B Analytical derivations
We summarize the analytical derivation of the Holzapfel-Ogden model stress used in our code below:
$\bm{P}\left(\bm{F},\bm{\theta}\right)=$ $2\bm{F}\frac{\partial\psi(\bm{F})}{\partial\bm{C}}$ (32)
$=$ $2\bm{F}\left(\frac{\partial\bar{\psi}(\bar{I}_{1})}{\partial\bar{I}_{1}}\frac{\partial\bar{I}_{1}}{\partial\bm{C}}+\frac{\partial\bar{\psi}(\bar{I}_{4f})}{\partial\bar{I}_{4f}}\frac{\partial\bar{I}_{4f}}{\partial\bm{C}}\right$.
$+\left.\frac{\partial\bar{\psi}(\bar{I}_{4s})}{\partial\bar{I}_{4s}}\frac{\partial\bar{I}_{4s}}{\partial\bm{C}}+\frac{\partial\bar{\psi}(\bar{I}_{4fs})}{\partial\bar{I}_{4fs}}\frac{\partial\bar{I}_{4fs}}{\partial\bm{C}}+K\frac{\bm{J}^{2}-1}{2}\bm{C}^{-1}\right)$
$\frac{\partial\bar{\psi}(\bar{I}_{1})}{\partial\bar{I}_{1}}$ $=\frac{a}{2}exp(b(\bar{I}_{1}-3))$ (33)
$\frac{\partial\bar{\psi}(\bar{I}_{4i})}{\partial\bar{I}_{4i}}$ $=a_{i}\langle\bar{I}_{4i}-1\rangle\exp(b_{i}\langle\bar{I}_{4i}-1\rangle^{2})$
$\frac{\partial\bar{\psi}(\bar{I}_{4fs})}{\partial\bar{I}_{4fs}}$ $=a_{fs}\bar{I}_{4fs}exp(b_{fs}\bar{I}_{4fs}^{2})$
$\frac{\partial\bar{I}_{1}}{\partial\bm{C}}$ $=J^{-2/3}\frac{\partial I_{1}}{\partial\bm{C}}+I_{1}\frac{\partial J^{-2/3}}{\partial\bm{C}}=\bm{J}^{-2/3}\bm{I}-\frac{1}{3}\bar{I}_{1}\bm{C}^{-1}$ (34)
$\frac{\partial\bar{I}_{4i}}{\partial\bm{C}}$ $=J^{-2/3}\frac{\partial I_{4i}}{\partial\bm{C}}+I_{4i}\frac{\partial J^{-2/3}}{\partial\bm{C}}=\bm{J}^{-2/3}(\bm{i}\otimes\bm{i})-\frac{1}{3}\bar{I}_{4i}\bm{C}^{-1}$
$\frac{\partial\bar{I}_{4fs}}{\partial\bm{C}}$ $=J^{-2/3}\frac{\partial I_{4fs}}{\partial\bm{C}}+I_{4fs}\frac{\partial J^{-2/3}}{\partial\bm{C}}=\frac{1}{2}\bm{J}^{-2/3}(\bm{f}\otimes\bm{s}+\bm{s}\otimes\bm{f})-\frac{1}{3}\bar{I}_{4fs}\bm{C}^{-1}$
And the Kullback-Leibler derivation:
$KL\left[\pi(\bm{\theta},V_{s},\sigma^{2}|\bm{b})||q(\bm{\theta},V_{s},\sigma^{2})\right]$ $=KL\left[\pi(\bm{\Phi}|\bm{b})||q(\bm{\Phi})\right]=\int q(\bm{\Phi})log\left[\frac{q(\bm{\Phi})}{\pi(\bm{\Phi}|\bm{b})}\right]d\bm{\Phi}$ (35)
$=\int q(\bm{\Phi})log\left[\frac{q(\bm{\Phi})\pi(\bm{b})}{\pi(\bm{\Phi},\bm{b})}\right]d\bm{\Phi}$
$=\int q(\bm{\Phi})log\left[\frac{q(\bm{\Phi})}{\pi(\bm{\Phi},\bm{b})}\right]d\bm{\Phi}+\int q(\bm{\Phi})log[\pi(\bm{b})]d\bm{\Phi}$
$=\int q(\bm{\Phi})log\left[\frac{q(\bm{\Phi})}{\pi(\bm{\Phi},\bm{b})}\right]d\bm{\Phi}+log[\pi(\bm{b})]$
$=-\int q(\bm{\Phi})log\left[\frac{\pi(\bm{b}|\bm{\Phi})\pi(\bm{\Phi})}{q(\bm{\Phi})}\right]d\bm{\Phi}+log[\pi(\bm{b})]$
$=-\int q(\bm{\Phi})\left(log\left[\frac{\pi(\bm{\Phi})}{q(\bm{\Phi})}\right]+log[\pi(\bm{b}|\bm{\Phi})]\right)d\bm{\Phi}+log[\pi(\bm{b})]$
$=-\int q(\bm{\Phi})log\left[\frac{\pi(\bm{\Phi})}{q(\bm{\Phi})}\right]d\bm{\Phi}-\int q(\bm{\Phi})log[\pi(\bm{b}|\bm{\Phi})]d\bm{\Phi}+log[\pi(\bm{b})]$
Appendix C Noise-free orthotropic constitutive parameter inference results
@@
Hyperparameters circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
$\alpha_{vs}$ $1.914\times 10^{1}$ $1.928\times 10^{1}$ $1.928\times 10^{1}$ $1.916\times 10^{1}$ $1.930\times 10^{1}$ $1.930\times 10^{1}$
@@
$\beta_{vs}$ $7.401\times 10^{7}$ $7.201\times 10^{7}$ $7.169\times 10^{7}$ $924\times 10^{7} 7.7$$20\times 10^{7} 7.7$$714\times 10^{7}$
@@
$\sigma_{af}$ $4.608\times 10^{-9}$ $2.851\times 10^{-8}$ $2.972\times 10^{-8}$ $5.507\times 10^{-9}$ $4.784\times 10^{-8}$ $2.798\times 10^{-8}$
$\sigma_{bf}$ $2.990\times 10^{-5}$ $2.072\times 10^{-4}$ $2.109\times 10^{-4}$ $3.661\times 10^{-5}$ $6.346\times 10^{-4}$ $2.022\times 10^{-4}$
$\sigma_{an}$ $514\times 10^{-9} 1.1$$100\times 10^{-7}$ $8.551\times 10^{-9}$ $1.380\times 10^{-9}$ $1.084\times 10^{-7}$ $8.144\times 10^{-9}$
$\sigma_{bn}$ $6.611\times 10^{-5}$ $3.699\times 10^{-2}$ $4.123\times 10^{-4}$ $7.526\times 10^{-5}$ $3.117\times 10^{-2}$ $3.880\times 10^{-4}$
$\sigma_{afs}$ $1.906\times 10^{0}$ $4.088\times 10^{-5}$ $4.207\times 10^{-5}$ $6.564\times 10^{-5}$ $4.102\times 10^{-6}$ $1.034\times 10^{-5}$
$\sigma_{bfs}$ $1.370\times 10^{1}$ $4.410\times 10^{0}$ $3.996\times 10^{0}$ $1.283\times 10^{1}$ $4.534\times 10^{-2}$ $2.728\times 10^{-1}$
@@
$e_{rel,a}$ $4.765\times 10^{-5}$ $1.864\times 10^{-5}$ $2.925\times 10^{-5}$ $1.832\times 10^{-5}$ $1.644\times 10^{-5}$ $2.593\times 10^{-5}$
$e_{rel,b}$ $2.298\times 10^{-6}$ $3.659\times 10^{-5}$ $3.854\times 10^{-5}$ $1.223\times 10^{-5}$ $2.795\times 10^{-5}$ $3.428\times 10^{-5}$
$e_{rel,af}$ $5.495\times 10^{-5}$ $8.243\times 10^{-5}$ $7.541\times 10^{-5}$ $4.184\times 10^{-5}$ $1.356\times 10^{-5}$ $8.322\times 10^{-5}$
$e_{rel,bf}$ $1.471\times 10^{-6}$ $4.349\times 10^{-5}$ $4.047\times 10^{-5}$ $9.769\times 10^{-6}$ $8.735\times 10^{-6}$ $3.724\times 10^{-5}$
$e_{rel,an}$ $9.655\times 10^{-5}$ $1.844\times 10^{-4}$ $3.895\times 10^{-5}$ $7.867\times 10^{-5}$ $2.017\times 10^{-4}$ $4.702\times 10^{-5}$
$e_{rel,bn}$ $1.290\times 10^{-5}$ $2.250\times 10^{-4}$ $5.078\times 10^{-5}$ $1.396\times 10^{-5}$ $4.682\times 10^{-4}$ $4.864\times 10^{-5}$
$e_{rel,afs}$ $1.417\times 10^{2}$ $8.050\times 10^{-3}$ $1.901\times 10^{-3}$ $1.023\times 10^{-2}$ $2.520\times 10^{-3}$ $1.137\times 10^{-3}$
$e_{rel,bfs}$ $9.201\times 10^{-1}$ $1.716\times 10^{-1}$ $3.743\times 10^{-1}$ $8.794\times 10^{-1}$ $1.535\times 10^{-3}$ $6.900\times 10^{-3}$
$e_{rel,K}$ $2.423\times 10^{-6}$ $2.932\times 10^{-5}$ $4.502\times 10^{-5}$ $4.837\times 10^{-6}$ $1.661\times 10^{-5}$ $4.184\times 10^{-5}$
Appendix D Noise-affected orthotropic constitutive parameter inference results
@@
Hyperparameters circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
$\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$
$\alpha_{vs}$ 1.922e+01 1.930e+01 1.928e+01 1.922e+01 1.918e+01 1.923e+01
@@
$\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$
mean:
@@
$\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-4}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$ $\sigma=1e^{-3}$
$e_{rel,a}$ 4.632e-04 4.835e-03 4.466e-03 -1.355e-03 -2.248e-01 -1.418e-01
$e_{rel,b}$ -1.793e-04 -8.132e-03 -2.026e-03 -4.671e-02 -1.738e-01 -5.960e-02
$e_{rel,af}$ -4.334e-03 -4.031e-02 -1.806e-02 -4.855e-02 -9.203e-01 -6.801e-02
$e_{rel,bf}$ 1.847e-03 3.258e-02 1.073e-02 3.515e-02 5.563e-01 5.470e-02
$e_{rel,an}$ 5.049e-03 -4.218e-02 -7.769e-03 -2.432e-02 1.000e+$0$ -9.956e-02
$e_{rel,bn}$ -1.119e-03 9.979e-02 2.589e-04 2.969e-02 9.946e-01 2.204e-02
$e_{rel,afs}$ 1.000e+$0$ 5.264e-01 9.225e-01 1.000e+$0$ 1.000e+$0$ 1.000e+$0$
$e_{rel,bfs}$ 9.069e-01 -5.964e-01 -3.369e+$0$ 9.118e-01 9.657e-01 9.974e-01
$e_{rel,K}$ -1.137e-04 4.845e-03 -1.068e-03 7.770e-03 2.997e-01 8.453e-02

### v2

@@ 7 Conclusion
We presented a physics-informed stochastic variational framework that recovers the full constitutive parameter set of a nonlinear orthotropic hyperelastic model from a single biaxial stretch experiment. By exploiting intrinsic microstructural heterogeneity, or introducing simple geometric heterogeneity when needed, the method identifies not only the dominant tensile parameters but also shear-coupling terms that are typically weakly excited in biaxial tension. Our results demonstrate that an effective balance between underlying microstructure and biaxial loading improves identifiability, and that this balance can be set through slicing and tissue harvesting. Circumferential–radial slabs and rotated cuts that place both fiber and sheet directions in the loading plane enable recovery of the full parameter set under noise-free conditions. Under realistic displacement measurement noise, the approach remains robust and yields stresses and strain energies in close agreement with the ground-truth mechanical response, while providing posterior uncertainty estimates for parameters and predicted stresses and strain energies. Adding geometric heterogeneity, a central occlusion for example, systematically improves convergence and tightens posterior distributions, particularly for shear-related parameters. These results indicate that slicing protocols and experimental design can be used deliberately to enhance observability in full-field inverse characterization. More broadly, the framework enables localized characterization of orthotropic nonlinear biological tissues from a single heterogeneous full-field experiment. The proposed framework therefore provides a practical route to single-shot, uncertainty-aware constitutive characterization that is compatible with emerging full-field experimental modalities. This reduces experimental tissue manipulation, improves reproducibility, and supports studies where available tissue is limited.
Acknowledgements
M.P. acknowledges support through the NWO Veni Talent Award 20058.
@@ Appendix A Tissue slab geometry and loading conditions
Figure 1: Synthetic data generation: training sample dimensions and loading conditions Tissue slab geometry and applied boundary conditions for the training samples. We subject all training samples to a biaxial loading protocol $\lambda_{1}:\lambda_{2}-(1:2),(1:1),(2:1)$ up to $15\%$ of the initial length of the sample. All samples are of size 10 $\times$ 10 $\times$ 1 mm3. The geometrically heterogeneous tissue samples have a circular occlusion in the center with a radius of $R=1$ mm.
Appendix B Noise-free orthotropic constitutive parameter inference results
@@
circ-long-hom circ-rad-hom rad-heli-hom circ-long-het circ-rad-het rad-heli-het
$\alpha_{v_{s}}$ $1.914\times 10^{1}$ $1.928\times 10^{1}$ $1.928\times 10^{1}$ $1.916\times 10^{1}$ $1.930\times 10^{1}$ $1.930\times 10^{1}$
@@
$\beta_{v_{s}}$ $7.401\times 10^{7}$ $7.201\times 10^{7}$ $7.169\times 10^{7}$ $924\times 10^{7} 7.7$$20\times 10^{7} 7.7$$714\times 10^{7}$
@@
$\sigma_{a_{f}}$ $4.608\times 10^{-9}$ $2.851\times 10^{-8}$ $2.972\times 10^{-8}$ $5.507\times 10^{-9}$ $4.784\times 10^{-8}$ $2.798\times 10^{-8}$
$\sigma_{b_{f}}$ $2.990\times 10^{-5}$ $2.072\times 10^{-4}$ $2.109\times 10^{-4}$ $3.661\times 10^{-5}$ $6.346\times 10^{-4}$ $2.022\times 10^{-4}$
$\sigma_{a_{n}}$ $514\times 10^{-9} 1.1$$100\times 10^{-7}$ $8.551\times 10^{-9}$ $1.380\times 10^{-9}$ $1.084\times 10^{-7}$ $8.144\times 10^{-9}$
$\sigma_{b_{n}}$ $6.611\times 10^{-5}$ $3.699\times 10^{-2}$ $4.123\times 10^{-4}$ $7.526\times 10^{-5}$ $3.117\times 10^{-2}$ $3.880\times 10^{-4}$
$\sigma_{a_{fs}}$ $1.906\times 10^{0}$ $4.088\times 10^{-5}$ $4.207\times 10^{-5}$ $6.564\times 10^{-5}$ $4.102\times 10^{-6}$ $1.034\times 10^{-5}$
$\sigma_{b_{fs}}$ $1.370\times 10^{1}$ $4.410\times 10^{0}$ $3.996\times 10^{0}$ $1.283\times 10^{1}$ $4.534\times 10^{-2}$ $2.728\times 10^{-1}$
@@
$\hat{e}_{a}$ $4.765\times 10^{-5}$ $1.864\times 10^{-5}$ $2.925\times 10^{-5}$ $1.832\times 10^{-5}$ $1.644\times 10^{-5}$ $2.593\times 10^{-5}$
$\hat{e}_{b}$ $2.298\times 10^{-6}$ $3.659\times 10^{-5}$ $3.854\times 10^{-5}$ $1.223\times 10^{-5}$ $2.795\times 10^{-5}$ $3.428\times 10^{-5}$
$\hat{e}_{a_{f}}$ $5.495\times 10^{-5}$ $8.243\times 10^{-5}$ $7.541\times 10^{-5}$ $4.184\times 10^{-5}$ $1.356\times 10^{-5}$ $8.322\times 10^{-5}$
$\hat{e}_{b_{f}}$ $1.471\times 10^{-6}$ $4.349\times 10^{-5}$ $4.047\times 10^{-5}$ $9.769\times 10^{-6}$ $8.735\times 10^{-6}$ $3.724\times 10^{-5}$
$\hat{e}_{a_{n}}$ $9.655\times 10^{-5}$ $1.844\times 10^{-4}$ $3.895\times 10^{-5}$ $7.867\times 10^{-5}$ $2.017\times 10^{-4}$ $4.702\times 10^{-5}$
$\hat{e}_{b_{n}}$ $1.290\times 10^{-5}$ $2.250\times 10^{-4}$ $5.078\times 10^{-5}$ $1.396\times 10^{-5}$ $4.682\times 10^{-4}$ $4.864\times 10^{-5}$
$\hat{e}_{a_{fs}}$ $1.417\times 10^{2}$ $8.050\times 10^{-3}$ $1.901\times 10^{-3}$ $1.023\times 10^{-2}$ $2.520\times 10^{-3}$ $1.137\times 10^{-3}$
$\hat{e}_{b_{fs}}$ $9.201\times 10^{-1}$ $1.716\times 10^{-1}$ $3.743\times 10^{-1}$ $8.794\times 10^{-1}$ $1.535\times 10^{-3}$ $6.900\times 10^{-3}$
$\hat{e}_{K}$ $2.423\times 10^{-6}$ $2.932\times 10^{-5}$ $4.502\times 10^{-5}$ $4.837\times 10^{-6}$ $1.661\times 10^{-5}$ $4.184\times 10^{-5}$
Appendix C Noise-affected orthotropic constitutive parameter inference results
This appendix contains the numerical results for the parameter fitting procedure when displacement noise is applied to the experiments. In Table 5, we show the optimized hyperparameters at the final epoch. The parameters themselves and their accompanying standard deviation are reported in Table 6. Finally, we also report the relative error between the estimated material model parameters and the ground-truth in Table 7.
@@ Appendix C Noise-affected orthotropic constitutive parameter inference results
circ-long-het circ-rad-het rad-heli-het circ-long-het circ-rad-het rad-heli-het
$\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-3}$ $\sigma_{u}=1e^{-3}$ $\sigma_{u}=1e^{-3}$
$\alpha_{v_{s}}$ 1.922e+01 1.930e+01 1.928e+01 1.922e+01 1.918e+01 1.923e+01
@@
Parameters circ-long-het circ-rad-het rad-heli-het circ-long-het circ-rad-het rad-heli-het
$\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-3}$ $\sigma_{u}=1e^{-3}$ $\sigma_{u}=1e^{-3}$
Inferred $\mu_{\bm{\theta}}$:
@@
Parameters circ-long-het circ-rad-het rad-heli-het circ-long-het circ-rad-het rad-heli-het
$\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-4}$ $\sigma_{u}=1e^{-3}$ $\sigma_{u}=1e^{-3}$ $\sigma_{u}=1e^{-3}$
$\hat{e}_{a}$ 4.632e-04 4.835e-03 4.466e-03 -1.355e-03 -2.248e-01 -1.418e-01
$\hat{e}_{b}$ -1.793e-04 -8.132e-03 -2.026e-03 -4.671e-02 -1.738e-01 -5.960e-02
$\hat{e}_{a_{f}}$ -4.334e-03 -4.031e-02 -1.806e-02 -4.855e-02 -9.203e-01 -6.801e-02
$\hat{e}_{b_{f}}$ 1.847e-03 3.258e-02 1.073e-02 3.515e-02 5.563e-01 5.470e-02
$\hat{e}_{a_{n}}$ 5.049e-03 -4.218e-02 -7.769e-03 -2.432e-02 1.000e+$0$ -9.956e-02
$\hat{e}_{b_{n}}$ -1.119e-03 9.979e-02 2.589e-04 2.969e-02 9.946e-01 2.204e-02
$\hat{e}_{a_{fs}}$ 1.000e+$0$ 5.264e-01 9.225e-01 1.000e+$0$ 1.000e+$0$ 1.000e+$0$
$\hat{e}_{b_{fs}}$ 9.069e-01 -5.964e-01 -3.369e+$0$ 9.118e-01 9.657e-01 9.974e-01
$\hat{e}_{K}$ -1.137e-04 4.845e-03 -1.068e-03 7.770e-03 2.997e-01 8.453e-02
Appendix D Uni-modal deformation mode validation protocols and stress expressions
This appendix summarizes the homogeneous deformation modes used for validation in Section 5.1. We reproduce incompressible biaxial tension and triaxial shear protocols following [10, 11] and evaluate the resulting Cauchy stress responses.
We prescribe a spatially homogeneous deformation gradient
For biaxial tension, $\lambda_{\rm f}$ and $\lambda_{\rm n}$ are prescribed, incompressibility implies $\lambda_{\rm s}=1/(\lambda_{\rm f}\lambda_{\rm n})$, and all shear components are set to zero. The hydrostatic pressure is obtained by enforcing $\sigma_{\rm ss}=0$, which yields the following non-zero stresses
[…]
Here $\psi_{\star}$ denotes $\psi_{\star}=\partial\psi/\partial I_{\star}$.
For the triaxial simple shear cases we consider, $\lambda_{\rm f}=\lambda_{\rm s}=\lambda_{\rm n}\equiv 1$ and a single shear component is activated per test while all others are set to zero. For $\gamma_{\star}\geq 0$, the non-zero shear stresses read
Appendix E Mixed-modal invariant sampling and admissibility filtering
We sample mixed-modal deformation states by drawing invariants uniformly between the extrema induced by the homogeneous deformation modes in Eq. (31). We then reconstruct $\bm{C}$ in the local fiber-sheet-normal basis, compute $\sqrt{\bm{C}}$ via Schur decomposition, and evaluate the maximum principal Cauchy stress used for admissibility filtering.
We take the minimum and maximum values of the invariants induced by Eq. (31) and uniformly sample in between:
$I_{\star}\sim\mathcal{U}(I_{\star}^{\min},I_{\star}^{\max}),\quad I_{\star}\in\{I_{1},I_{4ff},I_{4nn},I_{4fs}\}.$ (34)
where $I_{\star}^{\min}$ and $I_{\star}^{\max}$ denote the minimum and maximum deformation invariants seen in the uni-modal biaxial tension and triaxial simple shear experiments studied in Section 5.1. We reconstruct the right Cauchy-Green tensor using the identity $I_{1}=I_{4ff}+I_{4ss}+I_{4nn}$ and assuming the same orientation basis as in the uni-modal spatially homogeneous loading case:
$C_{11}=I_{4ff},\qquad$ $C_{13}=C_{31}=C_{23}=C_{32}=0,$ (35)
$C_{22}=I_{1}-I_{4ff}-I_{4nn},\qquad$ $C_{12}=C_{21}=I_{4fs},$
$C_{33}=I_{4nn}.$
Using the Schur decomposition [85], we compute the right stretch tensor $\bm{U}=\sqrt{\bm{C}}$ and use the polar decomposition $\bm{F}=\bm{R}\bm{U}$. We then evaluate the Cauchy stress and extract the maximum principal stress:
$\sigma_{1}$ $=\mathcal{F}_{1}(J^{-1}\bm{F}\bm{S}\bm{F}^{T})=\mathcal{F}_{1}(J^{-1}\bm{R}\bm{U}\bm{S}\bm{U}\bm{R}^{T})=\mathcal{F}_{1}(J^{-1}\bm{U}\bm{S}\bm{U}),$ (36)
where $\mathcal{F}_{1}(\bm{\sigma})=\max\operatorname{eig}(\bm{\sigma})$ and $\bm{S}=\partial\Psi/\partial\bm{C}$. We accept only physiologically relevant samples satisfying $0.9999\leq J\leq 1.0001$, aligned with the tissue’s assumed quasi-incompressibility, and $\sigma_{1}\leq 20.3~\mathrm{kPa}$, aligned with the maximum principal stresses measured in experimental protocols on human myocardial tissue [11] (see Table 3 showing peak Cauchy biaxial tensile stresses in the left ventricle). We repeat sampling until 5,0005{,}000 states are retained.